OpenStudy (anonymous):

The two balanced equations (1) and (2) are for reactions in which gaseous carbon dioxide is produced from the combustion of (1) solid carbon and (2) gaseous carbon monoxide. 1. C(s) + O2(g) CO2(g) + 94.0 kcal 2. CO(g) + ½O2(g) CO2(g) + 67.6 kcal When 112 grams of carbon monoxide are consumed according to equation (2), which of the following occurs? (atomic weights: C = 12.0 g/mol, O = 16.0 g/mol)

thomaster (thomaster):

The question says: which of the following occurs? Where are the following?

OpenStudy (anonymous):

1.0 mole of carbon dioxide is produced. 67.6 kcal of heat are generated. 2.00 moles of oxygen are consumed. 0.25 mole of carbon dioxide is produced. 0.50 mole of oxygen is consumed. Here is the following

thomaster (thomaster):

Oke so we have to use reaction 2 \(\Large\sf CO(g)+\frac{1}{2}O_2(g)\to CO_2(g)\) Let's get rid of the fractionby multiplying the whole equation with 2. \(\Large\sf 2CO(g)+O_2(g)\to 2CO_2(g)\) 112 grams of CO is used in the reaction Convert that to moles. Do you know how to do that?

thomaster (thomaster):

Are you there @Miguel27525 ?

OpenStudy (anonymous):

No in dont know how...

thomaster (thomaster):

You divide 112 by the molar mass of carbon monoxide

OpenStudy (anonymous):

Umm so the molar mass of carbon monoxide is 28.01 right?

thomaster (thomaster):

C = 12.011 O = 15.999 CO=12.011+15.999=28.01 g/mol So yea :)

OpenStudy (anonymous):

So you divide 112/28.01?

thomaster (thomaster):

Yes, then you know how many moles of CO you have for the reaction

OpenStudy (anonymous):

0.25 moles?

thomaster (thomaster):

28.01 g/mol means 1 mol = 28.01 gram You have 112 gram, how can your answer be less than 1 mol? :P

OpenStudy (anonymous):

sorry sorry sorry :P it's 2.00 right?

thomaster (thomaster):

\(\large\dfrac{112}{28.01}=3.998~mol\)

OpenStudy (anonymous):

Ohhh

thomaster (thomaster):

for 2 mol CO, we get 2 mol CO2 So when we have 3.998 mol CO, we get 3.998 mol CO2 Only thing you have to do is convert 3.998 mol CO2 to grams. Multiply with the molar mass of CO2

OpenStudy (anonymous):

Thank you :D

thomaster (thomaster):

You're welcome :)